OpenStudy (anonymous):

A barn silo, excluding the top, is a circular cylinder. The silo is 10 m in diameter and the height is 25 m. Find the volume.

OpenStudy (aravindg):

use the formula of volume of cylinder volume =pi*r^2*h

OpenStudy (anonymous):

so it will be 3.14 ( 10 x 10) 25?

OpenStudy (anonymous):

@AravindG

Directrix (directrix):

volume =pi*r^2*h V = pi * 5 ^2 * 25 ---> @WrittenOnMyHeart Use the radius of the base rather than the diameter.

OpenStudy (anonymous):

Oh! Okay...

OpenStudy (anonymous):

I think I have it...

OpenStudy (aravindg):

V = pi * 5 ^2 * 25

OpenStudy (aravindg):

d=10 imples r=5